### v2

@@ Abstract
Do social networks and peer influence shape major life decisions in highly polarized settings? We explore this question by examining how peers influenced the allegiances of West Point cadets during the American Civil War. Leveraging quasi-random variations in the proportion of cadets from Free States, we analyze how these differences affected cadets’ decisions about which army to join. We have four main findings. First, there was a strong and significant peer effect: a higher proportion of classmates from Free States significantly increased the likelihood that cadets from Slave States joined the Union Army. Second, the peer effect interacted with geography and economic circumstances: almost all cadets from Free States joined the Union Army (if they decided to join the war), and most cadets from Slave States that had more than a third of their population in slavery joined the Confederacy. The cadets who were most highly influenced were from Slave States that had less than a third of their population in slavery. Third, we analyze how cadets’ decisions affected their military rank and career outcomes. Fourth, we show that having served together in the Mexican-American war increased the peer influence, providing additional evidence that peer interaction was influential.
@@ 1 Introduction
Do social networks and peer influence shape individuals’ choices when societies are deeply polarized? This question is central not only to understanding moments of historical rupture but also to how political allegiances are formed in contemporary democracies. The literature on polarization has highlighted partisan sorting, media biases, divergent economic situations, and cultural divisions as key drivers of political alignment (e.g., Mason, 2018; Prior, 2013; Autor et al., 2020; Hochschild, 2016). Less is known about the role of peers and social networks in shaping political loyalties in polarized settings. While research has shown that peers matter for outcomes such as education, consumption, health, and careers, [^1] We cannot include a comprehensive review of the literature here, but such overviews are available in Benhabib, Bisin and Jackson (2010), Brechwald and Prinstein (2011), Aral and Nicolaides (2017), McGloin and Thomas (2019), Giletta et al. (2021), and Boucher et al. (2024). there is little causal evidence as to whether peers influence high-stakes political and/or life decisions, especially when these decisions come with severe personal risks and moral conflicts.
@@
The American Civil War provides a unique opportunity to examine this question. West Point cadets decisions as to which side to fight for was often the most consequential choice of their lives, pitting personal backgrounds, economic interests, and regional loyalties against professional ties and peer influences. Although West Pointers played pivotal roles in the war, systematic empirical analysis of the factors influencing their allegiances remains limited. [^2] There exists a rich historical literature on West Pointers in the Civil War (e.g., Patterson, 2002). A recent political science study, White (2024), highlights the importance of economic interests. See also additional references below. This setting offers two key advantages. First, the quasi-random composition of cadet cohorts generates exogenous variation in exposure to peers from different regions. This allows us to separate peer effects from confounding factors (e.g., individuals typically self-select into like-minded groups) that typically make it difficult to disentangle network influence from shared beliefs or communal forces (see Manski, 1993; Angrist, 2014). Quasi-randomness arose from institutional features: although admissions broadly tracked the geographic distribution of congressional districts, annual recruitment varied substantially due to decentralized nomination procedures and qualification exams (e.g., Park, 1840). Second, West Point was the nation’s premier military academy, where cadets from Northern (Free) and Southern (Slave) lived and trained together, and forged close bonds before the war. This environment magnified the potential for peer influence. Historians have long noted that “group cohesion” and camaraderie motivated enlistment and fighting, but the evidence has been largely anecdotal, drawn from letters and diaries (e.g., McPherson, 1997, 2003; Siebold, 2007). Whether Northern peers swayed Southern cadets toward the Union, or whether local political and economic pressures outweighed such influence, has been an open question.
@@
The first is that there is significant and large peer effect on cadet’s decisions. A higher proportion of classmates from Free States significantly increased the likelihood that cadets from Slave States joined the Union Army. The effect is sizable: a one-standard deviation increase in the proportion of peers from Free States (an increase of roughly 5 out of 40 peers) raised the probability of joining the Union by four percentage points, or 11% of the mean. In contrast, nearly all cadets from Free States joined the Union regardless of cohort composition. The difference suggests a conflict between nationalism and sectionalism faced by cadets from Slave States, who were more susceptible to peer influence.
The second main finding is that the peer effect depends on background (economic) situations. Within the Slave States, we find that peer influence mattered little for cadets from states in which the slave population was high (e.g., above 1/3), but mattered strongly for cadets from Slave States with slave populations below 1/3, suggesting an interaction between peer influence and economic considerations rooted in the slave economy. Indeed, the peer influence was large and almost entirely concentrated among cadets from Border and Mid-level Slave States where support for secession was more divided.
Our third main finding is that the cadets’ decisions had long-term implications for their careers and lives. Because many West Point cadets in our study later emerged as significant military figures, we can examine how their wartime choices influenced their post-war trajectories. Employing both OLS and an instrumental variable strategy (where home state and peer exposure predict allegiance), we find that for those from Slave States, those who sided with the Union experienced lower military ranks in 1865 and a diminished chance of becoming generals, but also a decreased likelihood of wartime mortality.
@@
Our fourth finding examines the role of direct interactions in driving peer influence. We present two pieces of evidence. We find that peer influence is more pronounced within the same cohort than across cohorts, indicating the importance of direct, contemporaneous interactions. We also leverage variation in cadets’ likelihood of having co-fought in the Mexican-American War (1846–1848). We find that peer effects were stronger when slave-state cadets had more free-state peers with whom they had shared combat experience, and weaker when their co-combatants were also from Slave States.
We present several additional analyses to examine heterogeneity and test the robustness of our results, including tests that account for shifting political environments, peer dropouts, those not participating in the war, appointment states, and other factors.
@@
Our study contributes knowledge about peer effects (see Benhabib, Bisin and Jackson (2010); Brechwald and Prinstein (2011); McGloin and Thomas (2019); Giletta et al. (2021) for some overviews), by demonstrating how peer influence can shape major life decisions, particularly during moments of extreme polarization. More broadly, it adds to the long-standing literature emphasizing the importance of “personal influence” in economic, political, and social behaviors (e.g., Katz and Lazarsfeld, 1955). Other research specifically involving soldiers concerns the spread of political ideas Jha and Wilkinson (2023); Grosjean et al. (2025), which our study complements in studying allegiance (and how it depends on economic circumstances).
@@
The experiences of West Point cadets during the American Civil War mirror other historical contexts where military academies produced leaders for both sides of major conflicts in divided nations. For instance, the Whampoa Military Academy in the 1920s produced commanders who fought on opposing sides during the Chinese Civil War (e.g., Bai, Jia and Wang, 2023), and the Heroic Military Academy in Mexico trained leaders who both supported and opposed the regime during the Mexican Revolution. Such decisions are not unique to military cadets, as similar choices are faced by ordinary citizens in many divided societies. However, analyzing peer influence in these broader contexts is often more challenging. The distinctive setting of the American Civil War, where allegiances were closely tied to geography, provides a unique opportunity to examine peer influence with a level of identification that is difficult to achieve in other cases.
@@
Our findings also add to the vast body of historical and social sciences literature that examines loyalty decisions during American Civil War. McPherson (1997) draws on soldiers’ letters and diaries to investigate soldiers’ motivations. Analyzing Union Army company data, Costa and Kahn (2003, 2008) emphasize the importance of loyalty to fellow soldiers, which sometimes surpassed commitment to the cause when considering decisions to fight, shirk duty, or desert. Our findings suggest that both cause (the allegiance to one’s home state) and comrades (peer influence) matter and that they also interact with each other. Closely related, Hall, Huff and Kuriwaki (2019) document that slave ownership drove enlistment in the Confederacy, whereas White (2024) explores the allegiance of West Pointers, emphasizing economic interest by showing that one’s past employment related to cash crops (i.e., sugar, indigo, rice, tobacco, and cotton) is linked to the likelihood of West Pointers joining the Confederacy. [^3] Using the same variable on cash-crop employment, we find a similar pattern. But with only a small share of candidates having such a history, we do not find a significant interaction effect between this variable and peer influence. Our observation regarding the influence of the slave economy corroborates these findings, though our primary focus remains on peer influence—a critical factor in loyalty decisions during pivotal historical moments.
@@
Finally, our study complements an expanding literature on leadership (e.g., Jones and Olken, 2005; Acemoglu and Jackson, 2015; Dippel and Heblich, 2021; Cagé et al., 2023; Bai, Jia and Yang, 2023; Ferrara, Dippel and Heblich, 2025). Understanding how leaders develop their ideas and loyalties is central to grasping important political and economic outcomes. Our findings highlight the significance of peer influence in shaping future leaders’ decisions during crises, demonstrating how peer influence can drive individual choices that may shape the course of history.
@@ 2.1 The Cadets
We focus on cadets who graduated from West Point Military Academy between 1820 and 1860. We manually collected information from the comprehensive Biographical Register of the Officers and Graduates of the United States Military Academy (Cullum, 1891). The records provide details such as the year of graduation, birth year, home state and appointment state, [^5] Home state and appointment state usually overlap. We use home state in our main analysis and present additional robustness checks using appointment state to define peers in the Supplemental Appendix. and academic rank. To ensure comparability of academic performance, we classify cadets into percentiles based on their graduation ranking. The top 10% were assigned a score of 100, the bottom 10% a score of 10, and so on. Additionally, the records include information on wars and battles one participated in and their military ranks, which we use to determine if one joined the Union or Confederate forces during the Civil War. We present examples of these records and our coding procedure in Supplemental Appendix A.1.
@@
Data on the economic conditions of cadets’ home states are sourced from the 1860 Census (United States Census Bureau, 1860). Our primary focus is on one key variable that serves as a proxy for economic conditions closely tied to the Civil War: the share of the slave population in the total population. Additionally, we examine other economic proxies, including logged farmland value per capita, logged manufactured product value per capita, and the share of manufacturing employment. These variables are typically highly correlated with each other and are absorbed by state fixed effects in our analysis. In additional analysis, we present the correlations between these variables and war allegiance, confirming that the share of the slave population is the most robust economic predictor of war allegiance.
@@ 2.2 Geography and the Classification of States by High/Low Slave
Between 1820 and 1860, 1,656 cadets graduated from the United States Military Academy (see the complete records in Biographical Register of the Officers and Graduates of the United States Military Academy (Cullum, 1891)). 17 graduates were from outside the continental United States, and since there was only one graduate in 1813, with no peers, our analytical sample consists of 1,638 cadets. In our data, 964 and 674 cadets came from Free States and Slave States, respectively. Among those from Free States, 540 joined the armies and 91.9% of them joined the Union. Among those from Slave States, 389 joined the armies and 36.2% of them joined the Union.
@@ 3.2 Statistical Analysis
where $Union_{i,s,t}$ denotes whether cadet $i$ from state $s$ and cohort $t$ joined the Union.
$Peer_{i,t}$ represents the proportion of peers from Free States. [^12] Note that our specification avoids the exclusion bias noted by Caeyers and Fafchamps (2024), since we are separating regressions for Slave and Free States and working with exogenous characteristics. $\alpha_{s}$ represents state fixed effects, whereas $X_{i,s}$ includes a range of personal and state characteristics such as the cadet’s age in 1860, their academic rank at West Point and the proportion of the slave population in their home state. Naturally, state characteristics are absorbed once home state fixed effects are accounted for.
We incorporate a linear time trend $t$ to account for potential temporal trends in Union enlistment. In addition, we report estimates that incorporate five-year bin fixed effects, which help absorb shifts in the political environment. To address potential serial correlation, we compute bootstrapped standard errors based on 400 resampling iterations. For robustness, we also re-estimate the model using standard errors clustered at the cohort level as well as bootstrapped standard errors clustered at the cohort level.
@@ 3.3 The Impact of Peers on Cadets’ Allegiance
Table 1 displays our main baseline results, indicating that peer influence matters for individuals from Slave States. Column (1) illustrates the raw correlation between the proportion of Free-State peers and the likelihood of joining the Union. Column (2) incorporates personal and state characteristics, while Column (3) further includes state fixed effects. To facilitate the interpretation of coefficients, we present the coefficient for a one standard deviation change in peer exposure. As shown, these estimates, between 0.040 and 0.062, are consistent across different models.
@@
The peer effect is sizable: a one standard deviation increase in of free state peers is roughly 5.5 more out of 40 cadets [^13] See Appendix Supplement Table S3. in the cohort being from Free States and leads to an 11 to 16% increase in the frequency of joining the union (mean 0.36). This is a conservative estimate. When using alternative thresholds for Slave States (5% and 10%), we find that peer effect raises the probability of Slave-State cadets joining the union by 17-19% (see Supplemental Appendix Table S6).
@@
(1) (2) (3) (4) (5) (6)
Share Free-State peers (sd) 0.062∗∗ 0.045∗∗ 0.040 0.006 0.004 0.002
@@
Age in 1860 -0.014 -0.002 0.007 0.005
(0.016) (0.016) (0.007) (0.007)
Class rank 0.001 0.001 -0.001 -0.001
(0.001) (0.001) (0.000) (0.000)
Slave pop. share (sd) -0.214∗∗∗ -0.009
@@
Cohort -0.014 -0.003 0.009 0.007
(0.015) (0.015) (0.006) (0.007)
State FEs N N Y N N Y
Dependent var. mean 0.362 0.362 0.362 0.919 0.919 0.919
Observations 389 389 389 540 540 540
R-squared 0.017 0.220 0.268 0.000 0.015 0.028
@@
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Columns (1)-(3) focus on cadets from Slave States (i.e., with a slave population higher than 1%) and Columns (4)-(6) on those from Free States (i.e., with a slave population lower than 1%). The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
@@
Among the personal background variables, the share of the slave population in a cadet’s home state is strongly and negatively associated with joining the Union, underscoring the role of economic factors in allegiance decisions. Supplemental Appendix Table S7 presents correlations between the probability of joining the Union and additional state-level economic proxies, including logged farmland value per capita, logged manufactured product value per capita, and the share of manufacturing employment. Notably, incorporating these proxies does not alter our primary findings on peer effects. Furthermore, the share of the slave population emerges as the most robust economic predictor of war allegiance. All these economic variables are effectively accounted for when controlling for state fixed effects.

### v3

@@ Abstract
Do social networks and peer influence shape major life decisions in highly polarized settings? We explore this question by examining how peers influenced the allegiances of West Point cadets during the American Civil War. Leveraging quasi-random variations in the proportion of cadets from Free States, we analyze how these differences affected cadets’ decisions about which army to join. We have three main findings. First, there was a strong and significant peer effect: a higher proportion of classmates from Free States significantly increased the likelihood that cadets from Slave States joined the Union Army. Second, the peer effect varies with geography, most notably with the slave population share in cadets’ home states or counties, and with cadets’ own slave ownership in 1860. Third, shared experiences—such as having served together in the Mexican-American War, continuous military service, and belonging to the same cohort—amplified peer effects, suggesting that sustained interaction is important.
@@ 1 Introduction
Do social networks and peer influence shape individuals’ choices when societies are deeply polarized? This question is central not only to understanding moments of historical rupture but also to how political allegiances are formed in contemporary democracies. The literature on polarization has highlighted partisan sorting, media biases, divergent economic situations, and cultural divisions as key drivers of political alignment (e.g., Mason, 2018; Prior, 2013; Autor et al., 2020; Hochschild, 2016). Less is known about the role of peers and social networks in shaping political loyalties in polarized settings. While research has shown that peers matter for outcomes such as education, consumption, health, and careers, [^1] We cannot include a comprehensive review of the literature here, but such overviews are available in Benhabib et al. (2010), Brechwald and Prinstein (2011), Aral and Nicolaides (2017), McGloin and Thomas (2019), Giletta et al. (2021), and Boucher et al. (2024). there is little causal evidence as to whether peers influence high-stakes political and/or life decisions, especially when these decisions come with severe personal risks and moral conflicts.
@@
The American Civil War provides a unique opportunity to examine this question. West Point cadets decisions as to which side to fight for was often the most consequential choice of their lives, pitting personal backgrounds, political-economic interests, and regional loyalties against professional ties and peer influences. Although West Pointers played pivotal roles in the war, systematic empirical analysis of the factors influencing their allegiances remains limited. [^2] There exists a rich historical literature on West Pointers in the Civil War (e.g., Patterson, 2002). A recent political science study, White (2024), highlights the importance of economic interests. See also additional references below. This setting offers two key advantages. First, the quasi-random composition of cadet cohorts generates exogenous variation in exposure to peers from different regions. This allows us to separate peer effects from confounding factors (e.g., individuals typically self-select into like-minded groups) that typically make it difficult to disentangle network influence from shared beliefs or communal forces (see Manski, 1993; Angrist, 2014). Quasi-randomness arose from institutional features: although admissions broadly tracked the geographic distribution of congressional districts, annual recruitment varied substantially due to decentralized nomination procedures and qualification exams (e.g., Park, 1840). Second, West Point was the nation’s premier military academy, where cadets from Northern (Free) and Southern (Slave) lived and trained together, and forged close bonds before the war. This environment magnified the potential for peer influence. Historians have long noted that “group cohesion” and camaraderie motivated enlistment and fighting, but the evidence has been largely anecdotal, drawn from letters and diaries (e.g., McPherson, 1997, 2003; Siebold, 2007). Whether Northern peers swayed Southern cadets toward the Union, or whether local political and economic pressures outweighed such influence, has been an open question.
@@
Our primary finding is that there is significant and large peer effect on cadet’s decisions. A higher proportion of classmates from Free States significantly increased the likelihood that cadets from Slave States joined the Union Army. The effect is sizable: a one-standard deviation increase in the proportion of peers from Free States (an increase of roughly 5 out of 40 peers) raised the probability of joining the Union by 5.4 percentage points, or 15% of the mean. In contrast, nearly all cadets from Free States joined the Union regardless of cohort composition. The difference suggests a conflict between nationalism and sectionalism faced by cadets from Slave States, who were more susceptible to peer influence. We formalize this intuition with a simple framework that illustrates why there are asymmetric peer effects between cadets from Slave and Free States.
The framework also implies that peer effects vary with political-economic background, which shapes the tradeoff inherent in the allegiance decision. Using slave-holding data from the 1860 Census and genealogical records from Ancestry.com and Familysearch.com, we find consistent evidence for this prediction. Among cadets from Slave States, stronger ties to slavery, proxied by a very high slave population shares in one’s home state (or county) or by personal slave ownership, essentially eradicate peer influence. The peer effect is very large in regions and cadets less tied to slavery. Moreover, slave population share emerges as a more powerful proxy for political-economic background than alternative measures, including county-level presence of pro- or anti-slavery religious groups and county-level voting shares for pro- and anti-slavery candidates in the 1860 presidential election.
@@
Our third set of analyses highlights the central role of interaction in shaping peer influence. We find that peer effects are substantially stronger within cohorts than across cohorts, indicating the importance of direct and contemporaneous interaction. Peer influence is also more pronounced among cadets who maintained continuous military careers between their graduation and the Civil War, compared with those who had returned to civilian life, suggesting that sustained institutional embeddedness amplified interpersonal influence. We also track variation in cadets’ likelihood of having served together in the Mexican–American War (1846–1848). Peer effects are stronger when cadets from Slave States had more Free-State peers with whom they shared combat experience, and weaker when their co-combatants were also from Slave States. Together, these findings provide additional evidence that peer influence operates through repeated/sustained interactions.
In our additional analysis, we examine heterogeneity and test the robustness of our results, including tests that account for shifting political environments, peer dropouts, those not participating in the war, appointment states, and other factors. We also trace some of implications that cadets’ decisions had on their careers and lives. Because many West Point cadets in our study later emerged as significant military figures, we can examine how their wartime choices influenced their post-war trajectories. Employing both OLS and an instrumental variable strategy (where home state and peer exposure predict allegiance), we find that for those from Slave States, those who sided with the Union experienced lower military ranks in 1865 and a diminished chance of becoming generals, but also a decreased likelihood of wartime mortality.
@@
Our study contributes knowledge about peer effects (see Benhabib et al. (2010); Brechwald and Prinstein (2011); McGloin and Thomas (2019); Giletta et al. (2021) for some overviews), by demonstrating how peer influence can shape major life decisions, particularly during moments of extreme polarization. More broadly, it adds to the long-standing literature emphasizing the importance of “personal influence” in economic, political, and social behaviors (e.g., Katz and Lazarsfeld, 1955). Other research specifically involving soldiers concerns the spread of political ideas (Jha and Wilkinson, 2023; Grosjean et al., 2025), which our study complements in studying allegiance (and how it depends on economic circumstances).
@@
The experiences of West Point cadets during the American Civil War mirror other historical contexts where military academies produced leaders for both sides of major conflicts in divided nations. For instance, the Whampoa Military Academy in the 1920s produced commanders who fought on opposing sides during the Chinese Civil War (e.g., Bai et al., 2023a), and the Heroic Military Academy in Mexico trained leaders who both supported and opposed the regime during the Mexican Revolution. Such decisions are not unique to military cadets, as similar choices are faced by ordinary citizens in many divided societies. However, analyzing peer influence in these broader contexts is often more challenging. The distinctive setting of the American Civil War, where allegiances were closely tied to geography, provides a unique opportunity to examine peer influence with a level of identification that is difficult to achieve in other cases.
@@
Our findings also add to the vast body of historical and social sciences literature that examines loyalty decisions during American Civil War. McPherson (1997) draws on soldiers’ letters and diaries to investigate soldiers’ motivations. Analyzing Union Army company data, Costa and Kahn (2003, 2008) emphasize the importance of loyalty to fellow soldiers, which sometimes surpassed commitment to the cause when considering decisions to fight, shirk duty, or desert. Our findings suggest that both cause (the allegiance to one’s home state) and comrades (peer influence) matter and that they also interact with each other. Closely related, Hall et al. (2019) document that slave ownership drove enlistment in the Confederacy, whereas White (2024) explores the allegiance of West Pointers, emphasizing economic interest by showing that one’s past employment related to cash crops (i.e., sugar, indigo, rice, tobacco, and cotton) is linked to the likelihood of West Pointers joining the Confederacy. [^3] Using the same variable on cash-crop employment, we find a similar pattern. But with only a small share of candidates having such a history, we do not find a significant interaction effect between this variable and peer influence. Our observation regarding the influence of the slave economy corroborates these findings, though our primary focus remains on peer influence—a critical factor in loyalty decisions during pivotal historical moments.
@@
Finally, our study complements an expanding literature on leadership (e.g., Jones and Olken, 2005; Acemoglu and Jackson, 2015; Dippel and Heblich, 2021; Cagé et al., 2023; Bai et al., 2023b; Ferrara et al., 2025). Understanding how leaders develop their ideas and loyalties is central to grasping important political and economic outcomes. Our findings highlight the significance of peer influence in shaping future leaders’ decisions during crises, demonstrating how peer influence can drive individual choices that may shape the course of history.
@@ 2.1 The Cadets
We focus on cadets who graduated from West Point Military Academy between 1820 and 1860.555We begin in 1820 because we observe that some cadets from the earlier cohorts participated in the Civil War. We also report additional results that restrict the sample to younger cohorts only. We manually collected information from the comprehensive Biographical Register of the Officers and Graduates of the United States Military Academy (Cullum, 1891). The records provide details such as the year of graduation, birth year, home state and appointment state, [^6] Home state and appointment state usually overlap. We use home state in our main analysis and present additional robustness checks using appointment state to define peers in the Supplemental Appendix. and academic rank. To ensure comparability of academic performance, we classify cadets into percentiles based on their graduation ranking. The top 10% were assigned a score of 100, the bottom 10% a score of 10, and so on. Additionally, the records include information on wars and battles one participated in and their military ranks, which we use to determine if one joined the Union or Confederate forces during the Civil War. We present examples of these records and our coding procedure in Supplemental Appendix A.1.
@@
We collected additional data on cadet backgrounds including (i) a cadet’s home county, which can be linked to county-level characteristics; (ii) Cadet slave ownership (whether the cadet personally owned slaves in 1860); (iii) whether a cadet’s father and mother originated from Free States; and (iv) whether a cadet’s wife originated from a Free State (with unmarried cadets coded as not having such a spouse). These data are obtained by Ancestry.com and FamilySearch.com, including linked data from the 1860 Census. We describe our data-linking procedure in Supplemental Appendix A.1.
To proxy cadets’ political–economic background, we use the slave population share in a cadet’s home state and home county, as well as whether the cadet personally owned slaves in 1860. The first two measures are drawn from the 1860 Census (United States Census Bureau, 1860), and cadet slave ownership is obtained from Ancestry.com. We also examine a broader set of economic proxies (including log farmland value per capita, log manufactured output per capita, and the share of employment in manufacturing) as well as religious and voting measures. Across specifications, a cadet’s region’s slave population and the cadet’s slave ownership emerge as the most robust predictors of wartime allegiance.
@@ 2.2 Geography and the Classification of States by High/Low Slave
Between 1820 and 1860, 1,656 cadets graduated from the United States Military Academy (see the complete records in Biographical Register of the Officers and Graduates of the United States Military Academy (Cullum, 1891)). 17 graduates were from outside the continental United States, and one was born in Indian Territory, our analytical sample consists of 1,638 cadets. In our data, 968 and 670 cadets came from Free States and Slave States, respectively. Among those from Free States, 540 joined the armies and 91.9% of them joined the Union. Among those from Slave States, 388 joined the armies and 36.1% of them joined the Union.
@@ 3.2 Statistical Analysis
where $Union_{i,s,t}$ denotes whether cadet $i$ from state $s$ and cohort $t$ joined the Union. $Peer_{i,t}$ represents the proportion of peers from Free States. [^13] Note that our specification avoids the exclusion bias noted by Caeyers and Fafchamps (2024), since we are separating regressions for Slave and Free States and working with exogenous characteristics. $\alpha_{s}$ represents state fixed effects.
$X_{i,s}$ includes a range of personal characteristics. Among them, our main analysis focuses on variables available for all cadets recorded in (Cullum, 1891): age in 1860, academic rank at West Point, and the slave population share in the cadet’s home state. When home-state fixed effects are included, state-level characteristics are naturally absorbed. In addition, we consider alternative measures of political-economic background, including whether a cadet owned slaves in 1860, and indicators for whether the cadet’s father, mother, or wife originated from Free States. These variables are constructed for the subset of cadets for whose information is available from Ancestry.com and FamilySearch.com.
We incorporate a linear time trend $t$ to account for potential temporal trends in Union enlistment. In addition, we report estimates that incorporate five-year bin fixed effects, which help absorb shifts in the political environment. To address potential serial correlation, we compute bootstrapped standard errors based on 400 resampling iterations.
Our specification examines peer effects based on exogenous peer characteristics and not endogenous peer choices (Manski (1993)). Given that cadets’ decisions were made some years after graduation, it is less likely that there was direct coordination on decisions rather than lasting impacts of interactions from their formative years. Nonetheless, it is possible that there is correlation in their decisions, and thus for robustness, we also re-estimate the model using standard errors clustered at the cohort level as well as bootstrapped standard errors clustered at the cohort level. We also directly examine residual dependence across peer groups by estimating intra-class correlation coefficients (ICC) for the baseline residuals within cohorts. The results show negligible within-cohort residual correlation, providing no evidence of jointly determined wartime allegiance decisions.
@@ 3.3 The Impact of Peers on Cadets’ Allegiance
Table 1 displays our main baseline results, indicating that peer influence matters for individuals from Slave States. Column (1) illustrates the raw correlation between the proportion of Free-State peers and the likelihood of joining the Union. Column (2) incorporates personal and state characteristics, while Column (3) further includes state fixed effects. To facilitate the interpretation of coefficients, we present the coefficient for a one standard deviation change in peer exposure. In Column (4), we consider the subset of cadets with additional individual background information. Across all specifications, we observe a strong peer influence for individuals from Slave States.
The peer effect is sizable: a one standard deviation increase in of free state peers is roughly 5.6 more out of 40 cadets [^14] See Appendix Supplement Table S3. in the cohort being from Free States and leads to an 15% increase in the frequency of joining the union (mean 0.36). This is a conservative estimate. When using alternative thresholds for Slave States (5% and 10%), we find that peer effect raises the probability of Slave-State cadets joining the union by 21% (see Supplemental Appendix Table S5). The effect is above 30% when we examine cadets who were not slave owners in 1860 (discussed later).
@@
(1) (2) (3) (4) (5) (6) (7) (8)
Share Free-State Peers (sd) 0.080∗∗∗ 0.058∗∗∗ 0.054∗∗∗ 0.056∗∗ -0.003 -0.005 -0.006 -0.005
(0.023) (0.020) (0.020) (0.022) (0.013) (0.013) (0.014) (0.015)
Age in 1860 -0.012 -0.005 0.004 0.006 0.004 -0.001
@@
Class Rank 0.001 0.001 0.002∗∗ -0.001∗∗ -0.001∗∗ -0.001∗∗∗
(0.001) (0.001) (0.001) (0.000) (0.000) (0.000)
Slave Pop. Share (sd) -0.216∗∗∗ -0.007
@@
Cohort -0.013 -0.005 0.002 0.008 0.007 0.003
(0.015) (0.015) (0.016) (0.007) (0.007) (0.008)
Cadet Slave Ownership -0.276∗∗∗
Free-State Father 0.090 0.019
(0.070) (0.061)
Free-State Mother 0.077 0.034
(0.075) (0.060)
Free-State Wife 0.061 0.098∗∗∗
(0.057) (0.028)
State FEs N N Y Y N N Y Y
Dependent var. mean 0.361 0.361 0.361 0.361 0.919 0.919 0.919 0.919
Observations 388 388 388 324 540 540 540 456
R-squared 0.028 0.234 0.287 0.363 0.000 0.017 0.030 0.067
@@
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Columns (1)-(4) focus on cadets from Slave States (i.e., with a slave population higher than 1%) and Columns (5)-(8) on those from Free States (i.e., with a slave population lower than 1%). The variable Free-State father and Free-State mother equal 1 if a parent is from a Free State, and 0 otherwise. A subset of observations contains missing information on parental origin. Free-State wife equals 1 if the wife is from a Free State and 0 otherwise. If the wife’s information is missing, we assign a value of 0. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
We include cadets who graduated between 1820 and 1860 in our baseline analysis because even some cadets from the early cohorts (from 1820) participated in the Civil War, often serving in officer or administrative roles. As a robustness check, we re-estimate our baseline specifications after excluding those aged 60 (or 50) and above; the estimated peer effects remain quantitatively similar and statistically robust. The results are reported in Appendix Table S7.
@@
Among the personal background variables, the share of the slave population in a cadet’s home state is strongly and negatively associated with joining the Union, underscoring the role of political-economic factors in allegiance decisions. Supplemental Appendix Table S8 presents correlations between the probability of joining the Union and additional state-level economic proxies, including logged farmland value per capita, logged manufactured product value per capita, and the share of manufacturing employment. Notably, incorporating these proxies does not alter our primary findings on peer effects. Furthermore, the share of the slave population emerges as the most robust economic predictor of war allegiance. All these economic variables are effectively accounted for when controlling for state fixed effects.
The individual-level cadet slave ownership is also a strong predictor of allegiance. As expected, it is negatively correlated with the probability of joining the Union.